# Uni-Instruct: One-step Diffusion Model through Unified Diffusion Divergence Instruction

### v2

@@ Abstract
In this paper, we unify more than 10 existing one-step diffusion distillation approaches, such as Diff-Instruct, DMD, SIM, SiD, $f$-distill, etc, inside a theory-driven framework which we name the Uni-Instruct. Uni-Instruct is motivated by our proposed diffusion expansion theory of the $f$-divergence family. Then we introduce key theories that overcome the intractability issue of the original expanded $f$-divergence, resulting in an equivalent yet tractable loss that effectively trains one-step diffusion models by minimizing the expanded $f$-divergence family. The novel unification introduced by Uni-Instruct not only offers new theoretical contributions that help understand existing approaches from a high-level perspective but also leads to state-of-the-art one-step diffusion generation performances. On the CIFAR10 generation benchmark, Uni-Instruct achieves record-breaking Frechet Inception Distance (FID) values of 1.46 for unconditional generation and 1.38 for conditional generation. On the ImageNet-$64\times 64$ generation benchmark, Uni-Instruct achieves a new SoTA one-step generation FID of 1.02, which outperforms its 79-step teacher diffusion with a significant improvement margin of 1.33 (1.02 vs 2.35). We also apply Uni-Instruct on broader tasks like text-to-3D generation. For text-to-3D generation, Uni-Instruct gives decent results, which slightly outperforms previous methods, such as SDS and VSD, in terms of both generation quality and diversity. Both the solid theoretical and empirical contributions of Uni-Instruct will potentially help future studies on one-step diffusion distillation and knowledge transferring of diffusion models.
footnotetext: Correspondence to luoweijian@xiaohongshu.com.
@@ 1 Introduction
Figure 1: Left: selected FID scores of different models on ImageNet-$64\times 64$ conditional generation benchmark. Right: Conception overview of Uni-Instruct. The Uni-Instruct unifies more than 10 existing diffusion distillation methods in a wide range of applications. Please check Table 6 for details.
@@
Currently, the mainstream of training the one-step diffusion model is through proper distillation approaches that minimize divergences between distributions of the one-step model and some teacher diffusion models. For instance, Diff-Instruct(luo2023diff, ) was probably the first work that introduced one-step diffusion models by minimizing the Kullback-Leibler divergence. DMD (yin2024one, ) improves the Diff-Instruct by introducing an additional regression loss. Score-identity Distillation (SiD) (zhou2024score, ) studies the one-step diffusion distillation by minimizing the Fisher divergence, but without the proof of gradient equivalence of the loss function. Later, Score Implicit Matching (SIM)(luo2025one, ) introduced a complete proof of the gradient equivalence of losses that minimizes the general score-based divergence family, including the Fisher divergence as a special case. f-distill (xu2025one, ) and SiDA (zhou2024adversarial, ) recently generalized the Diff-Instruct and the SiD to the integral $f$-divergence and auxiliary GAN losses, resulting in performance improvements on image generation benchmarks. Other approaches have also elaborated on the one-step diffusion models on a wide range of applications through the lens of divergence minimization (hoang2023swiftbrush, ; zhou2024adversarial, ; zhou2024long, ; yin2024improved, ; luo2024diff, ; luo2024diffstar, ; huang2024flow, ).
Provided that existing one-step diffusion models have achieved impressive performances, with some of them even outperforming their multi-step teacher diffusions, existing training approaches seem conceptually separated into two lines:
(1). Diff-Instruct(luo2023diff, ), and its variants like DMD(yin2024one, ), tackle the Integral Kullback-Leibler Divergence, while f-distill(xu2025one, ) unifies the IKL as a special case of Integral f-divergence. These KL and $f$-divergence based distillation approaches have the advantage of fast convergence, but suffer from mode-collapse issues and sub-optimal performances;
(2). Score Implicit Matching (SIM(luo2025one, )) proves a solid theoretical equivalence of score-based divergence minimization, which unifies the SiD(zhou2024score, ) and Fisher divergences as special cases. Though these general score-based divergences minimization has shown surprising generation performance, they may suffer from slow convergence issues and sub-optimal fidelity.
Till now, it seems that the KL-based and Score-based divergence minimization approaches are pretty parallel in theory. Therefore, we are strongly motivated to answer an interesting yet important research question:
  • Can we unify KL-based and Score-based approaches in a unified theoretical framework? If we can, would the unified approach lead to better one-step diffusion models?
@@
In this way, we are able to unify more than 10 existing diffusion distillation methods across a wide range of applications via our proposed Uni-Instruct. The methods that have been unified by Uni-Instruct include both KL-divergence-based methods (such as Diff-Instruct(luo2023diff, ), DMD(yin2024one, ), and $f$-distill(xu2025one, )) and general score-divergence-based methods (such as Score Implicit Matching (SIM(luo2025one, )), SiD(zhou2024score, ), and SiDA(zhou2024adversarial, )), as is shown in Table 6. Such a novel unification of existing one-step diffusion models marks the uniqueness of Uni-Instruct, which brings new perspectives in understanding and connecting different one-step diffusion models. Besides the solid theoretical contributions, Uni-Instruct also leads to new State-of-the-art one-step image generation performances on competitive image generation benchmarks: it achieved a record-breaking FID (Fréchet Inception Distance) (heusel2017gans, ) value of 1.02 on the ImageNet64×$\times 64$ 64 conditional generation task. This score outperforms its 79-step teacher diffusion with a significant improvement margin of 1.33 (1.02 vs 2.35). Uni-Instruct also leads to new state-of-the-art one-step FIDs of 1.46 on CIFAR10 conditional generation task and 1.36 on CIFAR10 unconditional generation task, significantly outperforming previous one-step models such as f-distill, SiDA, SIM, SiD, DMD, and Diff-Instruct. It also outperforms competitive few-step generative models, including consistency models (geng2024consistency, ; song2023consistency, ; lu2024simplifying, ), moment matching distillation models (salimans2024multistep, ), inductive models (zhou2025inductive, ), and many others (xie2024distillation, ).
Besides the one-step generation benchmark, we are inspired by DreamFusion (poole2022dreamfusion, ), ProlificDreamer (wang2023prolificdreamer, ), and Diff-Instruct (luo2023diff, )’s application on 3D experiments. In Section 5.3, we also successfully apply Uni-Instruct as a knowledge transferring approach for text-to-3D generation applications, resulting in robust, diverse, and high-fidelity 3D contents which are slightly better than ProlificDreamer in quality and diversity.
@@
  • New SoTA Practical Performances: Uni-Instruct achieved new state-of-the-art generation performances (measured in FID) on CIFAR10 (a one-step FID of 1.36) and ImageNet64×$\times 64$ 64 (a one-step FID of 1.02) benchmarks. We also successfully applied Uni-Instruct on the text-to-3D generation task, resulting in plausible and diverse 3D generation results.
@@ 3 Uni-Instruct: Unify One-step Distillation Methods in Theory
In this section, we introduce Uni-Instruct, a theory-driven family of approaches for the one-step distillation of score-based diffusion models. Uni-Instruct is able to unify more than 10 existing methods as special cases with proper weighting functions. It also leads to new state-of-the-art one-step generation performances on ImageNet64×$\times 64$ 64 and CIFAR10 generation benchmarks.
@@ 3.1 Diffusion Expansion of $f$-Divergence
Appropriate choices of the function $f(\cdot)$ lead to many widely-used divergences such as reverse-KL divergence (RKL), forward-KL divergence (FKL), Jeffrey-KL divergence (JKL), Jensen-Shannon divergence (JS), and Chi-Square divergence ($\chi^{2}$). We put more introductions in the appendix B.
@@
We give a complete proof with regularity analysis in Appendix A.1. For simplicity, we assume $\mathcal{D}_{f}(q_{T}\|p_{\theta,T})=0$ and ignore the last term in Eq. 3.2 in the following section. This fundamental expansion (Eq. 3.2) expands the static $f$-divergence in data space into an integral of divergences along the diffusion process. However, directly optimizing objective (3.2) is not tractable because we do not know the exact expressions of either the density $p_{\theta,t}$ or the score function $\bm{s}_{p_{\theta,t}}(\cdot)$ of the diffused one-step model’s distribution. To step towards a tractable objective, we derive the $\theta$ gradient of the expanded $f$-divergence (3.2) in Theorem 3.2.
@@ 3.2 Theories to Get Tractable Losses
$=$ $-\frac{1}{2}g^{2}(t)\frac{\partial}{\partial\theta}\Bigg{\{}\mathbb{E}_{p_{\theta,t}}\Bigg{[}\text{SG}\Bigg{(}\mathcal{C}_{1}\left(\frac{q_{t}}{p_{\theta,t}}\right)\Bigg{)}\left(\bm{s}_{q_{t}}(\bm{x}_{t})-\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})\right)\left(\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\right)\Bigg{]}\Bigg{\}}$
@@
$\mathcal{L}_{\mathrm{SIM}}=\mathbb{E}_{p_{\theta,t}}\left[\text{SG}\left(\mathcal{C}_{1}\left(r\right)\right)\left(\bm{s}_{q_{t}}(\bm{x}_{t})-\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})\right)\left(\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\right)\right],$
@@
where the weighting coefficients are determined by the $f$-divergence selection, we provide our completed proofs in Appendix A.2.
@@
Notice that the tractable loss function (3.5) requires the density ratio between the one-step model and teacher diffusion. For this, we train a GAN discriminator along the process, where the discriminator output serves as an estimator. This use of GAN discriminator is also widely applicable in other works like SiDA(zhou2024adversarial, ) and $f$-distill (xu2025one, ). Details on why the GAN discriminator recovers the density ratio can be found in Theorem A.1.

### v4

@@ Abstract
In this paper, we unify more than 10 existing one-step diffusion distillation approaches, such as Diff-Instruct, DMD, SIM, SiD, $f$-distill, etc, inside a theory-driven framework which we name the Uni-Instruct. Uni-Instruct is motivated by our proposed diffusion expansion theory of the $f$-divergence family. Then we introduce key theories that overcome the intractability issue of the original expanded $f$-divergence, resulting in an equivalent yet tractable loss that effectively trains one-step diffusion models by minimizing the expanded $f$-divergence family. The novel unification introduced by Uni-Instruct not only offers new theoretical contributions that help understand existing approaches from a high-level perspective but also leads to state-of-the-art one-step diffusion generation performances. On the CIFAR10 generation benchmark, Uni-Instruct achieves record-breaking Frechet Inception Distance (FID) values of 1.46 for unconditional generation and 1.38 for conditional generation. On the ImageNet $64\times 64$ generation benchmark, Uni-Instruct achieves a new SoTA one-step generation FID of 1.02, which outperforms its 79-step teacher diffusion with a significant improvement margin of 1.33 (1.02 vs 2.35). We also apply Uni-Instruct on broader tasks like text-to-3D generation, which slightly outperform previous methods, such as SDS and VSD, in terms of both generation quality and diversity. Both the solid theoretical and empirical contributions of Uni-Instruct will potentially help future studies on one-step diffusion distillation and knowledge transfer of diffusion models. Code will be available at Github.
footnotetext: Correspondence to pkulwj1994@icloud.com.
@@ 1 Introduction
Figure 1: Left: Conception overview of Uni-Instruct. The Uni-Instruct unifies more than 10 existing diffusion distillation methods in a wide range of applications. Please check Table 1 for details. Right: selected FID scores of different models on ImageNet-$64\times 64$ conditional generation benchmark.
@@
Currently, the mainstream of training the one-step diffusion model is through proper distillation approaches that minimize divergences between distributions of the one-step model and some teacher diffusion models. For instance, Diff-Instruct(luo2023diff, ) was the first work that introduced one-step diffusion models by minimizing the Kullback-Leibler divergence. DMD (yin2024one, ) improves the Diff-Instruct by introducing an additional regression loss. Score-identity Distillation (SiD) (zhou2024score, ) studies the one-step diffusion distillation by minimizing the Fisher divergence, but without the proof of gradient equivalence of the loss function. Later, Score Implicit Matching (SIM)(luo2025one, ) introduced a complete proof of the gradient equivalence of losses that minimizes the general score-based divergence family, including the Fisher divergence as a special case. f-distill (xu2025one, ) and SiDA (zhou2024adversarial, ) recently generalized the Diff-Instruct and the SiD to the integral $f$-divergence and auxiliary GAN losses, resulting in performance improvements on image generation benchmarks. Other approaches have also elaborated on the one-step diffusion models on a wide range of applications through the lens of divergence minimization (hoang2023swiftbrush, ; zhou2024adversarial, ; zhou2024long, ; yin2024improved, ; luo2024diff, ; luo2024diffstar, ; huang2024flow, ).
Provided that existing one-step diffusion models have achieved impressive performances, with some of them even outperforming their multi-step teacher diffusions, existing training approaches can be conceptually separated into two lines:
(1) Diff-Instruct(luo2023diff, ), and its variants like DMD(yin2024one, ), tackle the Integral Kullback-Leibler Divergence, while f-distill(xu2025one, ) unifies the IKL as a special case of Integral f-divergence. These KL and $f$-divergence based distillation approaches have the advantage of fast convergence, but suffer from mode-collapse issues and sub-optimal performances;
(2) Score Implicit Matching (SIM(luo2025one, )) proves a solid theoretical equivalence of score-based divergence minimization, which unifies the SiD(zhou2024score, ) and Fisher divergences as special cases. Though these general score-based divergences minimization has shown surprising generation performance, they may suffer from slow convergence issues and sub-optimal fidelity.
Till now, it seems that the KL-based and Score-based divergence minimization approaches are pretty parallel in theory. Therefore, we are strongly motivated to answer an interesting yet important research question: Can we unify KL-based and Score-based approaches in a unified theoretical framework? If we can, would the unified approach lead to better one-step diffusion models?
@@
In this way, we are able to unify more than 10 existing diffusion distillation methods across a wide range of applications via our proposed Uni-Instruct. The methods that have been unified by Uni-Instruct include both KL-divergence based methods (such as Diff-Instruct(luo2023diff, ), DMD(yin2024one, ), and $f$-distill(xu2025one, )) and general score-divergence based methods (such as Score Implicit Matching (SIM(luo2025one, )), SiD(zhou2024score, ), and SiDA(zhou2024adversarial, )), as is shown in Table 1. Such a novel unification of existing one-step diffusion models marks the uniqueness of Uni-Instruct, which brings new perspectives in understanding and connecting different one-step diffusion models. Besides the solid theoretical contributions, Uni-Instruct also leads to new State-of-the-art one-step image generation performances on competitive image generation benchmarks: it achieved a record-breaking FID (Fréchet Inception Distance) (heusel2017gans, ) value of 1.02 on the ImageNet $64\times 64$ conditional generation task. This score outperforms its 79-step teacher diffusion with a significant improvement margin of 1.33 (1.02 vs 2.35). Uni-Instruct also leads to new state-of-the-art one-step FIDs of 1.46 on CIFAR10 conditional generation task and 1.36 on CIFAR10 unconditional generation task, significantly outperforming previous one-step models such as f-distill, SiDA, SIM, SiD, DMD, and Diff-Instruct. It also outperforms competitive few-step generative models, including consistency models (geng2024consistency, ; song2023consistency, ; lu2024simplifying, ), moment matching distillation models (salimans2024multistep, ), inductive models (zhou2025inductive, ), and many others (xie2024distillation, ).
Besides the one-step generation benchmark, we are inspired by DreamFusion (poole2022dreamfusion, ), ProlificDreamer (wang2023prolificdreamer, ), SIM (luo2025one, )’s application on 3D experiments. In Section 4.3, we also successfully apply Uni-Instruct as a knowledge transferring approach for text-to-3D generation applications, resulting in robust, diverse, and high-fidelity 3D contents which are slightly better than ProlificDreamer in quality and diversity.
@@
[…]
Method Loss Div. in UI Task Loss Function Gradient Expression
Diff-Instruct (DI) (luo2023diff, ) IKL Div. $\chi^{2}$ One-Step Diffusion $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $(\bm{s}_{p_{\theta,t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t}))\frac{\partial\mathbf{x}_{t}}{\partial\theta}$
DI++ (luo2024diff, ) IKL Div. + Reward $\chi^{2}$ Human Aligned One-Step Diffusion $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\mathcal{L}_{\text{reward}}$ $\text{Grad}({\mathrm{DI}})+\nabla_{\theta}\mathcal{L}_{\text{reward}}$
DI (luo2024diffstar, ) KL Div. + Reward RKL Human Aligned One-Step Diffusion $\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})$ $\mathcal{L}_{\text{reward}}$ $\text{Grad}(\mathrm{SIM})+\nabla_{\theta}\mathcal{L}_{\text{reward}}$
SDS (poole2022dreamfusion, ) IKL Div. $\chi^{2}$ Text to 3D $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\text{Grad}({\mathrm{DI}})$
DDS (hertz2023delta, ) IKL Div. $\chi^{2}$ Image Editing $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\text{Grad}({\mathrm{DI}})$
VSD (wang2023prolificdreamer, ) IKL Div. $\chi^{2}$ Text to 3D $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\text{Grad}({\mathrm{DI}})$
DMD (yin2024one, ) IKL Div. + Reg. $\chi^{2}$ One-Step Diffusion $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\mathcal{L}_{\text{MSE}}$ $\text{Grad}({\mathrm{DI}})+\nabla_{\theta}\text{MSE}$
RedDiff (mardani2023variational, ) IKL Div. + Data Fedility $\chi^{2}$ Inverse Problem $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\mathcal{L}_{\text{MSE}}$ $\text{Grad}({\mathrm{DI}})+\nabla_{\theta}\text{MSE}$
DMD2 (yin2024improved, ) IKL Div. + GAN $\chi^{2}$ One-Step Diffusion $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\mathcal{L}_{adv.}$ $\text{Grad}({\mathrm{DI}})+\nabla_{\theta}\mathcal{L}_{adv.}$
Swift Brush (hoang2023swiftbrush, ) IKL Div. $\chi^{2}$ One-Step Diffusion $\int w(t)\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})\mathrm{d}t$ $\text{Grad}({\mathrm{DI}})$
SIM (luo2025one, ) General KL Div. RKL One-Step Diffusion $\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})$ $\frac{\partial}{\partial\theta}\left(\bm{s}_{q_{t}}(\bm{x}_{t})-\bm{s}_{p_{\theta,t}}(\bm{x}_{t})\right)\cdot$ $\left(\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})-\nabla\log q_{t}(\bm{x}_{t}|\bm{x}_{0})\right)$
SiD (zhou2024score, ) KL Div. RKL One-Step Diffusion $\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})$ $\text{Grad}(\mathrm{SIM})$
SiDA (zhou2024adversarial, ) KL Div. + GAN RKL One-Step Diffusion $\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})$$\mathcal{L}_{adv.}$ $\text{Grad}({\mathrm{SIM}})+\nabla_{\theta}\mathcal{L}_{adv.}$
SiD-LSG (zhou2024long, ) KL Div. RKL One-Step Diffusion $\mathcal{D}_{\text{KL}}(p_{\theta,t}||q_{0})$ $\text{Grad}({\mathrm{SIM}})$
$f$-distill (xu2025one, ) I-$f$ Div. + GAN $\chi^{2}$ One-Step Diffusion $\int w(t)\mathcal{D}_{f}(q_{0}||p_{\theta,t})\mathrm{d}t$ $\mathcal{L}_{adv.}$ $\lambda_{f}\text{Grad}({\mathrm{DI}})+\nabla_{\theta}\mathcal{L}_{adv.}$
Uni-Instruct (Ours) $f$ Div. + GAN All All $\mathcal{D}_{f}(q_{0}||p_{\theta,t})$ $\mathcal{L}_{adv.}$ $\nabla_{\theta}\mathcal{L}_{adv.}+\lambda^{\mathrm{DI}}_{f}\text{Grad}({\mathrm{DI}})$ $\lambda^{\mathrm{SIM}}_{f}(\mathbf{x})\text{Grad}({\mathrm{SIM}})$
Table 1: Distribution matching diffusion distillation loss family. Uni-Instruct not only extends the distribution matching framework theoretically, but also unifies all previous gradient expressions with specific weightings.
@@ 3 Uni-Instruct: Unify One-step Distillation Methods in Theory
[…]
@@ 3.1 Diffusion Expansion of $f$-Divergence
Appropriate choices of the function $f(\cdot)$ lead to many widely-used divergences such as reverse-KL divergence (RKL), forward-KL divergence (FKL), Jeffrey-KL divergence (JKL), Jensen-Shannon divergence (JS), and Chi-Square divergence ($\chi^{2}$). We put more introductions in the appendix C.
@@
We give a complete proof with regularity analysis in Appendix B.1. For simplicity, we assume $\mathcal{D}_{f}(q_{T}\|p_{\theta,T})=0$ and ignore the last term in Eq. 3.2 in the following section. This fundamental expansion (Eq. 3.2) expands the static $f$-divergence in data space into an integral of divergences along the diffusion process. However, directly optimizing objective (3.2) is not tractable because we do not know the exact expressions of either the density $p_{\theta,t}$ or the score function $\bm{s}_{p_{\theta,t}}(\cdot)$ of the diffused one-step model’s distribution. To step towards a tractable objective, we derive the $\theta$ gradient of the expanded $f$-divergence (3.2) in Theorem 3.2.
@@ 3.2 Theories to Get Tractable Losses
$=$ $-\frac{1}{2}g^{2}(t)\frac{\partial}{\partial\theta}\Bigg\{\mathbb{E}_{p_{\theta,t}}\Bigg[\text{SG}\Bigg(\mathcal{C}_{1}\left(\frac{q_{t}}{p_{\theta,t}}\right)\Bigg)2\left(\bm{s}_{q_{t}}(\bm{x}_{t})-\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})\right)\left(\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\right)\Bigg]\Bigg\}$
@@
$\mathcal{L}_{\mathrm{SIM}}=\mathbb{E}_{p_{\theta,t}}\left[\text{SG}\left(\mathcal{C}_{1}\left(r\right)\right)2\left(\bm{s}_{q_{t}}(\bm{x}_{t})-\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})\right)\left(\bm{s}_{p_{sg[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\right)\right],$
@@
where the weighting coefficients are determined by the $f$-divergence selection, we provide our completed proofs in Appendix B.2.
@@
Notice that the tractable loss function (3.5) requires the density ratio between the one-step model and teacher diffusion. For this, we train a GAN discriminator along the process, where the discriminator output serves as an estimator. This use of GAN discriminator is also widely applicable in other works like SiDA(zhou2024adversarial, ) and $f$-distill (xu2025one, ). Details on why the GAN discriminator recovers the density ratio can be found in Theorem B.1.
@@
1:while not converge do
3:  Generate fake images: $\bm{x}_{\text{fake}}=g_{\theta}(\epsilon)$
5:  Update $s_{\psi}$ with denoising score matching loss:
[…]
9:end while
10:return $g_{\theta}$